OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = one thirdx + 14 thirds y = 3x + 10 y = -3x - 2 y = -one thirdx + 10 thirds

OpenStudy (anonymous):

Please help. :|

OpenStudy (stefrheart):

okay so first you need to put into point slope form

OpenStudy (anonymous):

Okay. Is that something like y2-y1/x2-x1? I forget. :|

OpenStudy (stefrheart):

no its the y-4=-3(x+2)

OpenStudy (anonymous):

Okay.

OpenStudy (stefrheart):

so then you need to distribute the -3

OpenStudy (anonymous):

Okay, so i distribute them through (x+2)?

OpenStudy (stefrheart):

yes

OpenStudy (anonymous):

I got -3x+1

OpenStudy (anonymous):

Well Y-4=-3x+1

OpenStudy (stefrheart):

okay so now you add the 4 on both sides

OpenStudy (anonymous):

Right, Thats because its passing through the equal sign, Correct?

OpenStudy (stefrheart):

well thats how you isolate the y

OpenStudy (anonymous):

Alright, So i ended up with Y=1x+1

OpenStudy (stefrheart):

wait you should have y=-3x+5

OpenStudy (anonymous):

Okay, That was my mistake on the math, Your right

OpenStudy (stefrheart):

oh wait hold on let me write it out i may be wrong

OpenStudy (anonymous):

Okay. I got it because i added 4 to -3

OpenStudy (stefrheart):

im sorry its y=-3x-2

OpenStudy (anonymous):

Okay Now what

OpenStudy (anonymous):

Oh wait, Thats the answer

OpenStudy (anonymous):

Thank you!

OpenStudy (stefrheart):

welcome!!